Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability Question: How to tackle this faster?

Expert replies
by Kaunteya » Wed Feb 20, 2008 11:52 am
I have issues with probability questions. I can get to the answer but it takes me a good 3-4 minutes. Can someone please explain to me an easier method to solve this question that may save me time. I have to always draw out a picture of the people ect...wasting valuable time.

In a room filled with 7 people, 4 people have exactly 1 friend in the room and 3 people have exactly 2 friends in the room (Assuming that friendship is a mutual relationship, i.e. if John is Peter's friend, Peter is John's friend). If two individuals are selected from the room at random, what is the probability that those two individuals are NOT friends?
a. 5/21
b. 3/7
c. 4/7
e. 5/7
f. 16/21

Merci beaucoup!
K
Join the discussion
Source: — Problem Solving |

by siddarthd2919 » Wed Feb 20, 2008 3:23 pm
can u post the answer.....if possible the steps.....i am stuck in this one.
Join the discussion

by Kaunteya » Wed Feb 20, 2008 3:41 pm
Begin by counting the number of relationships that exist among the 7 individuals whom we will call A, B, C, D, E, F, and G.

First consider the relationships of individual A: AB, AC, AD, AE, AF, AG = 6 total. Then consider the relationships of individual B without counting the relationship AB that was already counted before: BC, BD, BE, BF, BG = 5 total. Continuing this pattern, we can see that C will add an additional 4 relationships, D will add an additional 3 relationships, E will add an additional 2 relationships, and F will add 1 additional relationship. Thus, there are a total of 6 + 5 + 4 + 3 + 2 + 1 = 21 total relationships between the 7 individuals. Alternatively, this can be computed formulaically as choosing a group of 2 from 7: 7!/(2! 5!) = 21

We are told that 4 people have exactly 1 friend. This would account for 2 "friendship" relationships (e.g. AB and CD). We are also told that 3 people have exactly 2 friends. This would account for another 3 "friendship" relationships (e.g. EF, EG, and FG). Thus, there are 5 total "friendship" relationships in the group.

The probability that any 2 individuals in the group are friends is 5/21. The probability that any 2 individuals in the group are NOT friends = 1 – 5/21 = 16/21. The correct answer is E.

To get to the answer in 2-2.5 minute range is certainly not easy. I had to think about this one for about 3-4 minutes before I tried to answer it. Even then that took me way too long. Maybe someone out there has a shortcut for solving these types of questions.

Cheers
K
Join the discussion

This is the only way..

by deltasquare » Thu Feb 21, 2008 11:37 pm
The way Kaunteya has done it is the best and the only way to arrive at an answer for such questions.
If we try to do it any faster than this, we are likely to make a mistake while solving.
Such problems are typical of Manhattan tests, but its unlikely that such questions will occur in GMAT test unless we've been scoring very highly (getting a lot of questions in 700+ difficulty level right)..
To make it faster we can try to eliminate obviously wrong options and guess among the likely ones.. Can't think of any other way.
Join the discussion

by iwillmakeit » Sat Mar 01, 2008 9:22 pm
p=1-(4/7*1/6+3/7*2/6)=16/21

4/7 - the probability to choose one of 4 people who have 1 friend.
1/6 - the probability to choose a friend of 6 remained people

3/7 - the probability to choose one of 3 people who have 2 friend.
2/6 - the probability to choose a friend of 6 remained people.
Join the discussion